OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5.

OpenStudy (redohawk):

-3(x + 3) = -3(x + 1) - 5 -3x - 9 = -3x - 1 - 5 x - 9 = -1 - 5 x - 9 = -6 x = 3 (I think it is right)